Alumium sulphide has 3.2 gm of sulphur then find weight of aluminium  present in sample??

Grade:12th pass

1 Answers

Khimraj
5 years ago
Aluminium sulphide = Al2S3
for 2mole of Al = 3 mole of S
2*27 gm of Al = 3*32 gm of S
then 1 gm of Al = 16/9 gm of S
then 3.2 fm of Al = 51.2/9 = 5.7 gm of S
Hope it clears.